Findax/min point for the following function:
f (x,y) = -x^2-y^2 + 6x + 8y - 21

Answer by stanbon(75887)

Find max/min point for the following function:
f (x,y) = -x^2-y^2 + 6x + 8y - 21
Complete the square on the x and on the y terms::
-[x^-6x+9]-[y^2-8y+16] = 21-9-16 = -4
-----
(x-3)^2/4 + (y-4)^2/4 = 1
------------------
Circle with center at (3,4) and radius = 2
------
Max point (3,4+2) = (3,6)
Min point (3,4-2) = (3,2)
--------------
Cheers,
Stan H.
-------------